# BARI-TH/25-772

### v1

@@ Abstract
We consider an abelian extension of the Standard Model (SM) comprising a new gauge group $U(1)^{\prime}$, with the neutral gauge boson $Z^{\prime}$ having flavour violating couplings to quarks and leptons. The fermion content is the same as in SM except for the addition of three right-handed neutrinos. The model, proposed in [1], describes the couplings of $Z^{\prime}$ to fermions in terms of three rational parameters $\epsilon_{1,2,3}$ that sum to zero imposing the cancellation of the gauge anomalies. Each $\epsilon_{i}$ is common to all fermions in a generation, a feature producing correlations among quark and lepton observables. We focus on $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transitions for the lepton flavour conserving $\ell_{1}=\ell_{2}$ and lepton flavour violating case $\ell_{1}\neq\ell_{2}$. Small deviations with respect to the SM predictions are found in the first case, which reflects a feature of the model where quark and lepton sectors prevent each other to manifest large discrepancies with respect to SM. We investigate the correlations with the flavour violating leptonic decays $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$. The experimental upper bound on the branching ratio of $\mu^{-}\to e^{-}\gamma$ constrains the range for the lepton flavour violating $B_{d,s}$ decays, which however are predicted to be within the reach of the planned new facilities.
@@ 1 Introduction
The flavour changing neutral current (FCNC) modes are of particular interest. In the SM they are forbidden at tree-level, as a consequence of the universality of the weak interactions and of the unitarity of the Cabibbo-Kobayashi-Maskawa matrix (CKM) rotating quark flavour eigenstates to mass eigenstates.  [^1] A complete discussion is in [2]. Among the tensions, there are indeed anomalies in observables related to the FCNC $b\to s$ transition. Moreover, there are difficulties concerning the value of some CKM matrix elements. Hints of violation of lepton flavour universality have also been spotted. A more detailed list of discrepancies includes [3, 4]:
@@
  • Tensions related to the CKM matrix elements $|V_{cb}|$ and $|V_{ub}|$. In both cases the value measured from exclusive decay modes (namely $B\to D^{*}\mu\bar{\nu}_{\mu}$ and $B\to\pi\mu\bar{\nu}_{\mu}$) is smaller than the value obtained from the inclusive $B\to X_{c,u}\mu\bar{\nu}_{\mu}$ modes [5]. This long standing issue is still unsolved despite the numerous proposed explanations. [^2] For $|V_{cb}|$ a solution based on $\mu-e$ universality violation has been proposed in [6]. In addition, there is the so-called Cabibbo-angle anomaly, a $\sim 3\sigma$ tension among the determinations of $|V_{us}|$, together with a deficit from unitarity in the first row of the CKM matrix [7].
@@
  • Tensions in observables related to the $b\to s$ transition. The branching ratios ${\cal B}(B\to K\mu^{+}\mu^{-})$ and ${\cal B}(B_{s}\to\phi\mu^{+}\mu^{-})$ are below the SM estimates. For the mode $B\to K^{*}\mu^{+}\mu^{-}$ a number of angular observables can be defined, which depend on $q^{2}$, the squared dilepton invariant mass [8]. For low $q^{2}$ some observables exceed the SM prediction by about $3\sigma$ [5].
@@
  • Anomalies in semileptonic decays induced by $b\to c\ell\bar{\nu}_{\ell}$ transition ($\ell=e,\,\mu,\,\tau$). The measured ratios of branching fractions $R(D^{(*)})={\cal B}(B\to D^{(*)}\tau\bar{\nu}_{\tau})/{\cal B}(B\to D^{(*)}\ell\bar{\nu}_{\ell})$, with $\ell=e,\,\mu$, exceed the SM result: The current average of BaBar, Belle and LHCb Collaborations measurements deviates by about $3\sigma$ from the SM prediction [5]. This issue challenges the SM property of lepton flavour universality.
@@
In addition to the above tensions, a discrepancy appeared in $(g-2)_{\mu}$, the anomalous magnetic moment of the muon, in particular with the Muon $g-2$ Collaboration measurements at Fermilab [9]. In this observable the main error in the SM prediction is related to one of the hadronic contributions, the hadronic-vacuum-polarisation contribution, a nonperturbative QCD effect. A lattice QCD calculation of this contribution by the BMW Collaboration [10] differs by about $2.2\sigma$ from the one obtained by a data-driven dispersive method employing the hadronic $e^{+}e^{-}$ cross section, used for the SM prediction in the White Paper [11]. In this respect, the measurement of $\sigma(e^{+}e^{-}\to\pi^{+}\pi^{-})$ by the CMD-3 experiment is also in tension with previous results [12]. An updated analysis accounting for both effects shows that the discrepancy is drastically reduced [13]. Nevertheless, non SM contributions to $(g-2)_{\mu}$ deserve to be scrutinized.
@@
A single above mentioned deviation is not large enough to unambiguously signal NP. In general, the precision analysis is a complex task, since most observables involve SM parameters and nonperturbative quantities whose uncertainty need to be reliably assessed, as demonstrated by $(g-2)_{\mu}$. Methods have been proposed to get rid, to a large extent, of the errors of some SM inputs in a class of rare $K$ and $B$ decays [14, 15]. Taken all together the tensions could hint a BSM scenario, which however is severely constrained by the unreasonable effectiveness of the Standard Model.
@@
A minimal extension of SM has been investigated in [1], based on the additional $U(1)^{\prime}$ gauge group and comprising right-handed neutrinos. The focus was on a solution found for the set of equations enabling the cancellation of the gauge anomalies induced by the new $U(1)^{\prime}$. In this solution the $U(1)^{\prime}$ charges of the fermions are generation dependent, and are expressed in terms of the SM hypercharges and of two additional rational parameters. One consequence of this assignment is a set of correlations arising among quark observables, among lepton observables and, remarkably, in the quark-lepton sectors together. Some correlations have been investigated in [1]. The conclusion was that, within the model, quark and lepton sectors mutually act to prevent large deviations from SM, justifying the small discrepancy with SM in observables where the tensions have shown up. However, the path to new physics is still viable: it relies on lepton flavour violating processes forbidden in SM, the observation of which would be a smoking gun for BSM physics. Such a kind of processes are the main topic of the present study. We show that, within the model (from now on denoted as the ABCD model) some charged lepton flavour violating modes can occur at non negligible rates. Moreover, correlations can be established, for example among some rare exclusive decays induced by the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition and some lepton decays.  [^3] For recent analyses of lepton flavour violation in $b\to s$ modes see [16, 17, 18, 19].
@@
The plan of the paper is as follows: In Sec. 2 we describe the main features of the ABCD model, summarising the discussion in [1]. In Sec. 3 we select the set of observables to study within the model, and identify the correlations among them to investigate. The results of the numerical analysis are presented in Sec. 4. The conclusions are in the last section. Some details are included in the appendices. Appendix A collects the definition of the $B\to K^{*}$ matrix elements of the relevant local operators in terms of form factors. Appendix B contains the expressions of the angular coefficient functions of the $B\to K^{*}(K\pi)\ell^{-}_{1}\ell^{+}_{2}$ fully differential decay distribution. Appendix C includes the parametrization and the values of the CKM and the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrices used in this study.
@@ 2 The ABCD model
The ABCD model is a minimal extension of the SM gauge group with an additional $\text{U}(1)^{\prime}$ gauge group [1]. A new neutral gauge boson $Z^{\prime}$ is predicted, with gauge coupling $g_{Z^{\prime}}$. The charge associated to the $\text{U}(1)^{\prime}$ symmetry is named $z$-hypercharge. Many NP models are based on similar simple extensions, each one with specific $z$-hypercharge assignments [20, 21, 22, 23]. From the experimental side, the assumption on the hypercharges determines the exclusion plots in the plane of the $Z^{\prime}$ production cross section versus the mass $M_{Z^{\prime}}$.   [^4] See, e.g., the review: B.A. Dobrescu and S. Willocq, ”$Z^{\prime}$-boson searches”, in [24].
$M_{Z^{\prime}}$ sets the NP scale, and it is assumed to be acquired after spontaneous breaking of the new symmetry. In the model it is not specified how the SSB happens, the only assumption is that it occurs at a much higher scale than the SM Higgs vacuum expectation value. The mixing with other neutral gauge bosons is neglected. [^5] At tree-level the mixing vanishes in models where the SM Higgs is a singlet under $\text{U}(1)^{\prime}$. Besides $Z^{\prime}$, the model comprises heavy right-handed neutrinos considered as Dirac particles.
The quark sector is the same as in the SM: $q_{L}^{i}$ are $\text{SU}(2)_{L}$ left-handed quark doublets, $u_{R}^{i},\,d_{R}^{i}$ the right-handed singlets, $i=\{1,2,3\}$ is a generation index. In the lepton sector, $\ell_{L}^{i}$ denote the left-handed lepton doublets, $\nu_{R}^{i}$ and $\,e_{R}^{i}$ the right-handed singlets. Before the electroweak SSB the $Z^{\prime}$ couplings to fermions are flavour conserving: for a generic fermion $\psi$ the interaction Lagrangian with $Z^{\prime}$ reads
@@
After the enlargement of the gauge group, the SM remains free of gauge anomalies imposing anomaly cancellation conditions. The $z$-hypercharges are required to satisfy a set of six anomaly cancellation equations (ACE) [25], suitably written defining
@@
Several studies have been devoted to the rational solutions of Eqs. (7)-(12) [26]. The solution proposed in the ABCD model consists in writing the $z$-hypercharge of each fermion $f_{i}$ in terms of the SM hypercharge $y_{f_{i}}=y_{f}$ for all $i=1,2,3$ (hence generation universal) plus a rational parameter $\epsilon_{i}$ which depends on the generation $i$ and is the same for all fermions in the generation:
@@
Since the SM hypercharges $y_{f}$ solve the anomaly equations, the condition for $\epsilon_{i}$ to solve Eqs. (7)-(12) is simply [1]
@@
The parametrization of the rotation matrices can be found in [1]. Appendix C contains the choice for the parameters of the CKM and PMNS matrices. After the fermion rotations, flavour violating $Z^{\prime}$ couplings are generated, as indicated below.
@@
    The generation index $i=d,\,s,\,b$ is used to make the notation more transparent. $V_{ui}$ and $V_{ci}$ are CKM matrix elements.
@@
    The real numbers ${\tilde{s}}_{12},\,{\tilde{s}}_{13},\,{\tilde{s}}_{23}$ and the phases $\phi_{K},\,\phi_{d},\,\phi_{s}$ appear in the parametrization of the rotation matrices [1]. $(\tilde{s}_{12},\,\phi_{K})$ only enter in the kaon sector, $(\tilde{s}_{13},\,\phi_{d})$ only in the $B_{d}$ sector, $(\tilde{s}_{23},\,\phi_{s})$ only in the $B_{s}$ sector.
@@
Since the number of parameters, even in this minimal extension of the Standard Model, is not small, a strategy for their treatment needs to be envisaged.
The general case where all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in Eqs.(36),(37) are taken to be different from zero was considered, allowing for $Z^{\prime}$ couplings to all right-handed fermions [1]. This possibility, denoted as the scenario C, appeared not efficiently constrained. 666 In particular, the efficient constraints imposed by the $\Delta F=2$ observables discussed in the following were not imposed. Another strategy was adopted in two other cases. First, all $\tilde{t}_{ij}$ in (37) were assumed to vanish, enabling flavour violating couplings of $Z^{\prime}$ to right-handed fermions only for quarks, a case denoted as scenario B. Finally, all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in (36),(37) were set to zero, enabling no flavour violating couplings of $Z^{\prime}$ to right-handed fermions [1]. In this last case, the scenario A, the new parameters are only those in (35) together with the elements of the CKM and PMNS matrices. This is the most economic scenario assumed in our study. In the following, we select a set of relevant observables, we discuss how the parameter space is constrained and the implications in a few lepton flavour conserving (LFC) and lepton flavour violating processes (LFV). Since all $Z^{\prime}$ couplings to fermions involve the two parameters $\epsilon_{1,2}$, correlations among different modes can be established: we investigate some of them, in particular the correlations between quark and lepton observables.

### v2

@@ Abstract
We consider an abelian extension of the Standard Model (SM) comprising a new gauge group $U(1)^{\prime}$, with the neutral gauge boson $Z^{\prime}$ having flavour violating couplings to quarks and leptons. The fermion content is the same as in SM except for the addition of three right-handed neutrinos. The model, proposed in [Aebischer:2019blw], describes the couplings of $Z^{\prime}$ to fermions in terms of three rational parameters $\epsilon_{1,2,3}$ that sum to zero imposing the cancellation of the gauge anomalies. Each $\epsilon_{i}$ is common to all fermions in a generation, a feature producing correlations among quark and lepton observables. We focus on $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transitions for the lepton flavour conserving $\ell_{1}=\ell_{2}$ and lepton flavour violating case $\ell_{1}\neq\ell_{2}$. Small deviations with respect to the SM predictions are found in the first case, which reflects a feature of the model where quark and lepton sectors prevent each other to manifest large discrepancies with respect to SM. We investigate the correlations between rare $B$ and $B_{s}$ decays and the leptonic processes $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}\gamma$, $\mu^{-}\to e^{-}e^{+}e^{-}$ and the $\mu^{-}\to e^{-}$ conversion in nuclei. We show that the current experimental upper bounds on these four channels play an increasingly important role in constraining the branching fractions of lepton flavour violating $B$ and $B_{s}$ decays. While the present bound on $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ does not impose significant restrictions, the other three modes set progressively more stringent limits, an important information for the planned new experimental facilities.
@@ 1 Introduction
The flavour changing neutral current (FCNC) modes are of particular interest. In the SM they are forbidden at tree-level, as a consequence of the universality of the weak interactions and of the unitarity of the Cabibbo-Kobayashi-Maskawa matrix (CKM) rotating quark flavour eigenstates to mass eigenstates.  [^1] A complete discussion is in [Buras:2020xsm]. Among the tensions, there are indeed anomalies in observables related to the FCNC $b\to s$ transition. Moreover, there are difficulties concerning the value of some CKM matrix elements. Hints of violation of lepton flavour universality have also been spotted. A more detailed list of discrepancies includes [Colangelo:2024ped, Capdevila:2023yhq]:
@@
  • Tensions related to the CKM matrix elements $|V_{cb}|$ and $|V_{ub}|$. In both cases the value measured from exclusive decay modes (namely $B\to D^{*}\mu\bar{\nu}_{\mu}$ and $B\to\pi\mu\bar{\nu}_{\mu}$) is smaller than the value obtained from the inclusive $B\to X_{c,u}\mu\bar{\nu}_{\mu}$ modes [HeavyFlavorAveragingGroupHFLAV:2024ctg]. This long standing issue is still unsolved despite the numerous proposed explanations. [^2] For $|V_{cb}|$ a solution based on $\mu-e$ universality violation has been proposed in [Colangelo:2016ymy]. In addition, there is the so-called Cabibbo-angle anomaly, a $\sim 3\sigma$ tension among the determinations of $|V_{us}|$, together with a deficit from unitarity in the first row of the CKM matrix [Kitahara:2024azt].
@@
  • Tensions in observables related to the $b\to s$ transition. The branching ratios ${\cal B}(B\to K\mu^{+}\mu^{-})$ and ${\cal B}(B_{s}\to\phi\mu^{+}\mu^{-})$ are below the SM estimates. For the mode $B\to K^{*}\mu^{+}\mu^{-}$ a number of angular observables can be defined, which depend on $q^{2}$, the squared dilepton invariant mass [Altmannshofer:2008dz]. For low $q^{2}$ some observables exceed the SM prediction by about $3\sigma$ [HeavyFlavorAveragingGroupHFLAV:2024ctg].
@@
  • Anomalies in semileptonic decays induced by $b\to c\ell\bar{\nu}_{\ell}$ transition ($\ell=e,\,\mu,\,\tau$). The measured ratios of branching fractions $R(D^{(*)})={\cal B}(B\to D^{(*)}\tau\bar{\nu}_{\tau})/{\cal B}(B\to D^{(*)}\ell\bar{\nu}_{\ell})$, with $\ell=e,\,\mu$, exceed the SM result: The current average of BaBar, Belle and LHCb Collaborations measurements deviates by about $3\sigma$ from the SM prediction [HeavyFlavorAveragingGroupHFLAV:2024ctg]. This issue challenges the SM property of lepton flavour universality.
@@
In addition to the above tensions, a discrepancy appeared in $(g-2)_{\mu}$, the anomalous magnetic moment of the muon, in particular with the measurements at BNL [Muong-2:2006rrc] and Fermilab [Muong-2:2021ojo, Muong-2:2023cdq]. In this observable the main uncertainty in the SM prediction is related to one of the hadronic contributions, the leading-order hadronic-vacuum-polarisation contribution (LO HVP), a nonperturbative QCD effect. A lattice QCD calculation of this contribution by the BMW Collaboration [Borsanyi:2020mff] differs by about $2.2\sigma$ from the one obtained by a data-driven dispersive method employing the hadronic $e^{+}e^{-}$ cross section, used for the SM prediction in the White Paper 2020 of the g2g-2 Theory Initiative [Aoyama:2020ynm]. In this respect, the measurement of $\sigma(e^{+}e^{-}\to\pi^{+}\pi^{-})$ by the CMD-3 experiment is also in tension with previous results [CMD-3:2023alj]. An analysis which excludes results by data-driven dispersive methods and only uses lattice QCD determinations of the LO HVP contribution shows that measurements and SM expectations agree within the errors, which are dominated by the theory uncertainties [Aliberti:2025beg]. The agreement persists for the full data set of the Muon g2g-2 Collaboration and for the combination with all previous measurements [Muong-2:2025xyk]. Nevertheless, non SM contributions to $(g-2)_{\mu}$ still deserve to be scrutinized.
@@
A single above mentioned deviation is not large enough to unambiguously signal NP. In general, the precision analysis is a complex task, since most observables involve SM parameters and nonperturbative quantities whose uncertainty need to be reliably assessed, as demonstrated by $(g-2)_{\mu}$. Methods have been proposed to get rid, to a large extent, of the errors of some SM inputs in a class of rare $K$ and $B$ decays [Buras:2021nns, Buras:2022wpw]. Taken all together the tensions could hint a BSM scenario, which however is severely constrained by the unreasonable effectiveness of the Standard Model.
@@
A minimal extension of SM has been investigated in [Aebischer:2019blw], based on the additional $U(1)^{\prime}$ gauge group and comprising right-handed neutrinos. The focus was on a solution found for the set of equations enabling the cancellation of the gauge anomalies induced by the new $U(1)^{\prime}$. In this solution the $U(1)^{\prime}$ charges of the fermions are generation dependent, and are expressed in terms of the SM hypercharges and of two additional rational parameters. One consequence of this assignment is a set of correlations arising among quark observables, among lepton observables and, remarkably, in the quark-lepton sectors together. Some correlations have been investigated in [Aebischer:2019blw]. The conclusion was that, within the model, quark and lepton sectors mutually act to prevent large deviations from SM, justifying the small discrepancy with SM in observables where the tensions have shown up. However, the path to new physics is still viable: it relies on lepton flavour violating processes forbidden in SM, the observation of which would be a smoking gun for BSM physics. Such a kind of processes are the main topic of the present study. Within the considered framework (from now on denoted as the ABCD model), we specifically investigate possible correlations among charged lepton flavour violating (LFV) processes. We first analyse the interplay between rare beauty meson decays that conserve lepton flavour (LFC) and those exhibiting LFV. Subsequently, we explore potential correlations between processes induced by the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition and purely leptonic LFV modes.  [^3] Recent analyses of lepton flavour violating $b\to s$ modes are in [Crivellin:2015mga, Crivellin:2015lwa, Crivellin:2015era, Cornella:2021sby, Bordone:2021usz, Panda:2024ygr, Becirevic:2024vwy].
@@
The plan of the paper is as follows: In Sec. 2 we describe the main features of the ABCD model, summarising the discussion in [Aebischer:2019blw]. In Sec. 3 we select the set of observables to study within the model, and identify the correlations among them to investigate. The results of the numerical analysis are presented in Sec. 4. The conclusions are in the last section. Some details needed in the main text are included in the appendices: Appendix A collects the definition of the $B\to K^{*}$ matrix elements of the relevant local operators in terms of form factors. Appendix B contains the expressions of the angular coefficient functions of the $B\to K^{*}(K\pi)\ell^{-}_{1}\ell^{+}_{2}$ fully differential decay distribution. Appendix C includes the parametrization and the values of the CKM and the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrices used in the study.
@@ 2 The ABCD model
The ABCD model is a minimal extension of the SM gauge group with an additional $\text{U}(1)^{\prime}$ gauge group [Aebischer:2019blw]. A new neutral gauge boson $Z^{\prime}$ is predicted, with gauge coupling $g_{Z^{\prime}}$. The charge associated to the $\text{U}(1)^{\prime}$ symmetry is named $z$-hypercharge. Many NP models are based on similar simple extensions, each one with specific $z$-hypercharge assignments [Leike:1998wr, Langacker:2000ju, Appelquist:2002mw, Rizzo:2006nw]. From the experimental side, the assumption on the hypercharges determines the exclusion plots in the plane of the $Z^{\prime}$ production cross section versus the mass $M_{Z^{\prime}}$.   [^4] See, e.g., the review: B.A. Dobrescu and S. Willocq, $Z^{\prime}$-boson searches”, in [Navas:PDG].
$M_{Z^{\prime}}$ sets the NP scale, and it is assumed to be acquired after spontaneous breaking of the new symmetry. In the model it is not specified how the SSB happens, the only assumption is that it occurs at a much higher scale than the SM Higgs vacuum expectation value. We neglect the mixing with other neutral gauge bosons. The mass mixing vanishes at three-level if the SM Higgs is a singlet under $U(1)^{\prime}$. As for the kinetic mixing, it can be dealt with by recasting the gauge bosons kinetic terms in canonical form redefining the new gauge coupling [Holdom:1985ag, Babu:1996vt, DelAguila:1996fw]. Mixing generated by fermion loops can be neglected for processes mediated by $Z^{\prime}$ at tree-level.
Besides $Z^{\prime}$, the model comprises heavy right-handed neutrinos considered as Dirac particles. The quark sector is the same as in the SM: $q_{L}^{i}$ are $\text{SU}(2)_{L}$ left-handed quark doublets, $u_{R}^{i},\,d_{R}^{i}$ the right-handed singlets, $i=\{1,2,3\}$ is a generation index. In the lepton sector, $\ell_{L}^{i}$ denote the left-handed lepton doublets, $\nu_{R}^{i}$ and $e_{R}^{i}$ the right-handed singlets. Before the electroweak SSB the $Z^{\prime}$ couplings to fermions are flavour conserving: for a generic fermion $\psi$ the interaction Lagrangian with $Z^{\prime}$ reads
@@
After the enlargement of the gauge group, the SM remains free of gauge anomalies imposing anomaly cancellation conditions. The $z$-hypercharges are required to satisfy a set of six anomaly cancellation equations (ACE) [Carena:2004xs], suitably written defining
@@
Several studies have been devoted to the rational solutions of Eqs. (7)-(12) [Allanach:2018vjg]. The solution proposed in the ABCD model consists in writing the $z$-hypercharge of each fermion $f_{i}$ in terms of the SM hypercharge $y_{f_{i}}=y_{f}$ for all $i=1,2,3$ (hence generation universal) plus a rational parameter $\epsilon_{i}$ which depends on the generation $i$ and is the same for all fermions in the generation:
@@
Since the SM hypercharges $y_{f}$ solve the anomaly equations, the condition for $\epsilon_{i}$ to solve Eqs. (7)-(12) is simply [Aebischer:2019blw]
@@
The parametrization of the rotation matrices can be found in [Aebischer:2019blw]. Appendix C contains the choice for the parameters of the CKM and PMNS matrices. After the fermion rotations, flavour violating $Z^{\prime}$ couplings are generated, as indicated below.
@@
    The generation index i=d,s,bi=d,\,s,\,b is used to make the notation more transparent. $V_{ui}$ and $V_{ci}$ are CKM matrix elements.
@@
    The real numbers ${\tilde{s}}_{12},\,{\tilde{s}}_{13},\,{\tilde{s}}_{23}$ and the phases $\phi_{K},\,\phi_{d},\,\phi_{s}$ appear in the parametrization of the rotation matrices [Aebischer:2019blw]. $(\tilde{s}_{12},\,\phi_{K})$ only enter in the kaon sector, $(\tilde{s}_{13},\,\phi_{d})$ only in the $B_{d}$ sector, $(\tilde{s}_{23},\,\phi_{s})$ only in the $B_{s}$ sector.
@@
Since the number of parameters, even in this minimal extension of the Standard Model, is not small, a strategy for their treatment needs to be envisaged. The general case where all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in Eqs.(36),(37) are taken to be different from zero was considered, allowing for $Z^{\prime}$ couplings to all right-handed fermions [Aebischer:2019blw]. This possibility, denoted as the scenario C, appeared not efficiently constrained.  [^5] In particular, the efficient constraints imposed by the $\Delta F=2$ observables discussed in the following were not imposed. Another strategy was adopted in two other cases. First, all $\tilde{t}_{ij}$ in (37) were assumed to vanish, enabling flavour violating couplings of $Z^{\prime}$ to right-handed fermions only for quarks, a case denoted as scenario B. Finally, all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in (36),(37) were set to zero, enabling no flavour violating couplings of $Z^{\prime}$ to right-handed fermions [Aebischer:2019blw]. In this last case, the scenario A, the new parameters are only those in (35) together with the elements of the CKM and PMNS matrices. This most economic scenario is assumed in our study. In the following, we select a set of relevant observables, we discuss how the parameter space is constrained and the implications in a few lepton flavour conserving (LFC) and lepton flavour violating processes (LFV). Since all $Z^{\prime}$ couplings to fermions involve the two parameters $\epsilon_{1,2}$, correlations among different modes can be established: we investigate some of them, in particular the correlations between quark and lepton observables.